OpenStudy (anonymous):

Write the series in summation notation: 100 + 95 + 90 + 85 + 80

jimthompson5910 (jim_thompson5910):

you are starting at 100 and decreasing by 5 each time so the basic nth term formula is an = a1 + d(n-1) an = 100 + (-5)(n-1) an = 100 - 5n + 5 an = -5n + 105

jimthompson5910 (jim_thompson5910):

then you are adding up the first 5 terms, which means you would write out